# AI Rules? Characterizing Reddit Community Policies Towards AI-Generated Content

### v2

@@ 3.1. The Broad Subreddit and Rules Subreddit Sets
In order to better understand how online communities across Reddit are responding to the arrival of AIGC, we conducted a large-scale data collection and analysis. In November 2024 we performed an automated crawl of Reddit’s “Top Communities” list [^1] https://www.reddit.com/best/communities/1/, which includes public subreddits with more than ten subscribers. We used a web-scraping library [^2] https://docs.scrapy.org/en/latest/ to paginate through this list and extract the IDs of all listed subreddits from the page HTML. We then used Reddit’s Data API [^3] https://support.reddithelp.com/hc/en-us/articles/16160319875092-Reddit-Data-API-Wiki to fetch metadata and rules for each subreddit ID.
From the $N=331,938$ subreddits that we saw in our crawl, we used a process summarized in Figure 1 to create several datasets for analysis. We used an automated tool [^4] https://pypi.org/project/langdetect/ to determine the language of each subreddit based on the text in its description and rules metadata. We removed subreddits with a primary language other than english or a creation date less than one month before the crawl date, in order to remove communities that had not yet had time to create rules. This yielded the Broad Subreddit Set of $N=307,543$ public, English-language subreddits that were at least one month old, which we used to understand platform-wide patterns in the data and factors that contribute to whether subreddits have rules at all.
Using this broad dataset, we constructed a second, smaller dataset: the Rules Subreddit Set ($N=99,969$). This dataset consists only of subreddits with non-empty rules metadata. We use this dataset to understand patterns specific to subreddits with rules. These two datasets are the focus of our descriptive analysis and were used to create two other datasets specific to AI rules, which are discussed in Sections 3.3.4 and 3.3.2.
@@ 3.2. Longitudinal Subreddit Set
To address RQ1’s concern with the shift in rule prevalence over time, we constructed a Longitudinal Subreddit Set, using a July 2023 data collection effort as a baseline, and identifying the subset of subreddits from the 2024 crawl that were seen in the earlier crawl. ChatGPT launched at the end of 2022, kicking off a period of increased public awareness and usage of AI. Given the dynamic nature of such a moment of flux, we decided to perform a longitudinal analysis on the same set of subreddits over a period of time, so that we could study the evolution of community norms and practices soon after this launch.
@@ 3.3. Data Labeling
To answer our research questions about rules governing the use of AI, we created multiple label categories for the AI rules and the communities that had rules. Through the process described below, we used a hybrid human-LLM coding technique to determine whether each of the subreddits in the Rules Subreddit Set and the Longitudinal Subreddit Set had an explicit rule about AI use, and to label them accordingly. We used these labels to identify the subset of subreddits with AI rules. We refer to the subset of the Rules Subreddit Set with AI rules as the AI Rules Subreddit Set ($N=4,251$). The AI-related rules from this set became our AI Rules Set ($N=4,458$). We used an iterative inductive coding technique to identify common types of AI rules. We then used a hybrid human-LLM coding technique to label the rules according to these categories.
Below, we describe the label categories we used to categorize subreddit AI rules (summarized in Table 2), as well as the human-LLM process that we used to label the data. We also describe how we determined and labeled the community type for each subreddit in the Rules Subreddit Set, according to several frameworks from existing literature (summarized in Table 2).
@@
3.3.1. Detecting Community Rules Governing AI Use
Community rules are often very long and complex, which made our datasets too large to manually code for the presence of rules governing the use of AI. To get a rough sense of the prevalence of these rules, we first used a simple regex to search the subreddits’ rules for AI-related keywords. This step found matches in the rules of less than $2\%$ of the subreddits. We then devised a protocol, inspired by recent HCI work using LLMs for qualitative labeling tasks (Choksi et al., 2024; Rao et al., 2024; Aubin Le Quéré et al., 2024), to detect AI-related rules across the larger datasets with more precision. The approach is based on comparing LLM-produced labels with those produced by human experts and iterating on the LLM prompt until desirable inter-rater reliability (IRR) is achieved between the human and LLM labels.
[…]
We used these manually labeled rules as ground truth against which to evaluate the performance of an LLM coder, which we planned to use to label the rest of the data. We crafted a prompt that combined the human coder instructions with Chain-of-Thought prompting techniques (Wei et al., 2023), including the recommended eight few-shot, in-context examples, which we selected from the manually labeled sample. We produced LLM labels by using OpenAI’s API to prompt gpt-4o-mini-2024-07-18, OpenAI’s current flagship small model, using this prompt and the same input data that were given to the human coders (a subreddit’s rules). We compared the LLM-produced labels with the human-produced ones and calculated an IRR of .7.7.7.7, which we deemed insufficient. Two human coders examined the discrepancies, discussed what they thought the correct label should be, and updated the prompt with a clarifying example. For example, one rule labeled NO by humans but YES by the LLM was about bots: “No bots(No bots): No bots. All bots are banned unless the bot provides exceptional value to /r/ MealPrepSunday.” To better align the LLM label with our preferred definition of an AI rule, we added this rule as a negative example in the prompt, with the explanation:“This rule bans bots, not AI. Bots are not necessarily AI.”
To improve IRR we repeated this process of randomly sampling, manually coding, generating LLM labels, comparing the results, and updating the prompt using the examples that the model got wrong. We performed three iterations of this step, at which point the LLM labels had an IRR of .95.95.95.95 with the human labels. We then labeled the entirety of both the Longitudinal Subreddit Set and the Rules Subreddit Set using OpenAI’s API and gpt-4o-mini-2024-07-18. The final prompt used can be found in Appendix A.1. This process identified the AI Rules Subreddit Set ($N=4,251$) of subreddits with rules governing AI use.
3.3.2. Identifying Emergent AI Rule Types
In order to better understand the types of rules that subreddits use to govern AI, we identified these rules in our sample and used an iterative inductive coding process to develop a descriptive taxonomy. To do this coding, we slightly modified the prompt developed and validated in Section 3.3.1 so that it would flag individual rules rather than a set of rules as being AI-related. Since the prompt required only a slight modification from the prior prompt, and since human coders would be manually reviewing the labels in the next step, we decided that we did not need to validate model performance before tagging the entire sample. We used this prompt to instruct gpt-4o-mini-2024-07-18 to label each of the $N=33,674$ individual rules in the AI Rules Subreddit Set, which produced the AI Rules Dataset ($N=4,458$) of individual rules about AI.
Next, two of the authors collaborated to identify emergent labels for the rules. The authors randomly selected $N=25$ rules from this dataset and independently performed affinity mapping (Harboe and Huang, 2015) by grouping rules that appeared similar into emergent categories. The authors then compared results and consolidated their categories into labels that best captured the elements of interest. Each of the two authors then independently applied this set of labels to another random sample of $N=25$ rules, once again grouping similar rules, including into new emergent groups where the existing labels did not suffice. The two authors repeated this process of randomly sampling, independently labeling, collaboratively comparing their results, and consolidating their labels, until no new labels emerged and they were confident that their taxonomy covered the relevant dimensions of variation in the dataset. The final label categories were Rationale, Stance, Medium and Framing, which are listed alongside some of their possible labels in Table 2 (see Table 6 for a full list of all labels and examples of tagged rules). We applied these labels to the entire AI Rules Dataset, as we describe next.
@@ 3.3.3. Applying Rule Classification Labels
We used LLMs to apply our emergent labels and labels from an additional Rule Type category from Fiesler et al. (Fiesler et al., 2018) to the entire AI Rules Dataset. The two first authors collaboratively developed a codebook with instructions for applying these labels. Fiesler et al. does not provide codebook instructions, so the authors interpreted the examples presented in the study and came up with instructions for how to apply each of the $24$ possible labels555“Rule Type” labels are not mutually exclusive, so we represented them as a series of binary labels, one for each possible Rule Type..
@@ 3.3.4. Applying Subreddit Classification Labels
We also sought to label the subreddits in the Rules Subreddit Set with details about the type of community that each subreddit was home to. To this end, we devised a process to apply labels from two label categories used in prior HCI literature: Topic labels from Fiesler et al. (Fiesler et al., 2018) and Community Archetype labels from Prinster et al. (Prinster et al., 2024). In applying labels, we considered a subreddit’s name, description, and rules metadata fields.
@@ 4. Findings
We group our findings by research question and present in sequence: (1) changes in the prevalence of rules governing AI use, (2) the types of communities that have these rules, and (3) the types of rules that we find for governing AI use.
4.1. RQ1: How Has the Prevalence of Subreddit Rules Governing the Use of AI Changed Over Time?
Table 3. Prevalence of rules and rules governing the use of AI for subreddits in the Longitudinal Subreddit Set. We show the prevalence at the time of each crawl as well as the % change between the crawls. In addition to the prevalence in the entire dataset, we include the prevalence in the top 1% largest subreddits, ranked by subscriber count at the time of the 2024 crawl.
Jul 2023 Nov 2024
N % N % % Change
@@ 4.1. RQ1: How Has the Prevalence of Subreddit Rules Governing the Use of AI Changed Over Time?
Table 3 shows the results of the detection task described in Section 3.3.1 on our Longitudinal Subreddit Set. We find that during the 2023 crawl, $N=1,348$ subreddits had rules governing the use of AI (.$6\%.$). By the time of our 2024 crawl, this number increased by $108.3\%$ to $N=2,808$ ($1.2\%$). Rules in general, and specifically rules governing the use of AI, are more common in larger subreddits. For example, $8.4\%$ of the largest $1\%$ of subreddits by subscriber count had AI rules during our first crawl. This percentage increased to $17.1\%$ in our second crawl. Figure 2 provides a more detailed demonstration of the differences in both the prevalence of AI rules and the increase in such rules among subreddits with different numbers of subscribers. The figure plots the number of subreddits in the Longitudinal Subreddit Set with AI rules at the time of each crawl. In order to expose the variation across subreddits of different sizes, we bucket subreddits along the X-axis into deciles based on their number of subscribers at the time of the second crawl (the labels represent the lower edge of the decile). Each decile contains approximately $22,773$ subreddits. For each decile we plot two bars representing the number of subreddits with AI rules: the left bar represents the number during the first crawl and the right bar represents the number during the second. Above the bars we include a numeric label representing the percent change between crawls. For example, the rightmost pair of bars shows that in the largest decile of subreddits, about $750$ had AI rules during the first crawl, and that this number increased by $128\%$ to about $1,750$ at the second crawl. The figure shows that subreddits with more subscribers are more likely to have AI rules, and that communities of all sizes saw increases in the prevalence of AI rules between our two crawls, though that increase was more prominent for larger communities.
For comparison, we additionally analyzed the prevalence of AI rules in subreddits in the Broad Subreddit Set that were created after our first crawl and were thus excluded from our Longitudinal Subreddit Set. The Broad Subreddit Set contained $N=30,962$ English language subreddits that were created after the first crawl (but were at least a month old during the second crawl). Of these newly created subreddits, $N=467$ had AI rules ($1.5\%$), higher than the percentage of subreddits with AI rules during the second crawl in the Longitudinal Subreddit Set ($1.2\%$). While this difference can be due to the fact that new subreddits have different characteristics, our separate analysis in the following section also suggests that newer subreddits are more likely to have AI rules.
Figure 2. Number of subreddits in the Longitudinal Subreddit Set with AI rules during each of our crawls. To demonstrate differences across subreddit size, subreddits are bucketed into deciles based on their subscriber count at the time of the second crawl. Each decile contains approximately $22,773$ subreddits. The bars shows the number of subreddits with AI rules in each decile and the labels above the bars indicate the percent change between crawls.
@@
The image contains a bar chart showing the number of AI-related rules in subreddits during two different crawls, with subreddits bucketed into deciles by subscriber count. The Y-axis represents the number of subeddits and the X-axis is subscriber count decile at the time of the second crawl (measured by the lower edge of each decile). Each decile has two bars above it: the left bar represents the 2023 crawl and the right represents the 2024 crawl. The number of subreddits with AI rules steadily increases as subscriber count grows. The largest decile (7,241) has the highest number, above 1750 during the secon crawl, while the smallest decile (10) has a number near 0.
@@ 4.2. RQ2: What Types of Communities Have Rules Governing AI Use?
In this section, we examine the characteristics of communities that have rules governing the use of AI. To do that, we compare the characteristics of two groups of subreddits in the Rules Subreddit Set: subreddits with rules about the use of AI, and those without such rules. We find that larger subreddits and those associated with Art or Celebrity topics are more likely to have explicit rules regulating AI.
Table 4. Differences between subreddits in the Rules Subreddit Set ($N=99,969$) with and without AI rules. For numerical subreddit characteristics, we show the average value across the subreddits in each group. For binary characteristics, we show the percentage of subreddits with this characteristic in each group and denote these percentages with (%). While we computed differences for all 28 topics, we show here only the ten topics with the largest absolute differences.
No AI Rule AI Rule Difference
Metadata Subscriber count 37,865.69 27,3571.19 235705.5***
Subreddit Age (Years) 4.94 6.32 1.38***
Subreddit created after ChatGPT (%) 24.78 31.33 6.56***
Community Archetype Content generation (%) 51.68 73.75 22.07***
Affiliation with an entity (%) 57.47 63.3 5.83***
Topical question and answer (%) 7.29 6.52 -0.78
Learning and perspective broadening (%) 20.52 16.61 -3.91***
Social support (%) 10.29 4.38 -5.91***
Topic Celebrity (%) 3.87 18.33 14.45***
Art (%) 2.36 12.99 10.63***
Entertainment (%) 19.01 22.96 3.95***
Politics (%) 2.21 0.54 -1.67***
Business and Organizations (%) 2.71 0.99 -1.73***
Games (%) 6.97 4.68 -2.29***
Local (%) 4.22 1.11 -3.11***
Hobby (%) 8.07 4.7 -3.36***
Humor (%) 9.75 3.32 -6.43***
Support (%) 8.98 2.52 -6.46***

### v3

@@ 3.1. The Broad Subreddit and Rules Subreddit Sets
In order to better understand how online communities across Reddit are responding to the arrival of AIGC, we conducted a large-scale data collection and analysis. In November 2024 we performed an automated crawl of Reddit’s “Top Communities” list [^2] https://www.reddit.com/best/communities/1/, which includes public subreddits with more than ten subscribers. We used a web-scraping library [^3] https://docs.scrapy.org/en/latest/ to paginate through this list and extract the IDs of all listed subreddits from each page. We then used Reddit’s Data API [^4] https://support.reddithelp.com/hc/en-us/articles/16160319875092-Reddit-Data-API-Wiki to fetch metadata and rules for each subreddit ID.
From the $N=331,938$ subreddits seen in our crawl, we used a process summarized in Figure 1 to create several datasets for analysis. We used an automated tool [^5] https://pypi.org/project/langdetect/ to determine the language of each subreddit based on the text in its description and rules metadata. We removed subreddits with a primary language other than English or a creation date less than one month before the crawl date, in order to remove communities that had not yet had time to create rules. This process yielded the Broad Subreddit Set of $N=307,543$ public, English-language subreddits that were at least one month old, which we used to understand platform-wide patterns in the data and factors that contribute to whether subreddits have rules at all.
Using this broad dataset, we constructed a second, smaller dataset: the Rules Subreddit Set ($N=99,969$). This dataset consists only of subreddits with non-empty rules metadata. We use this dataset to understand patterns specific to subreddits with rules. These two datasets are the focus of our descriptive analysis and were used to create two other datasets specific to AI rules, which are discussed in Sections 3.3.2 and 3.3.4.
@@ 3.2. Longitudinal Subreddit Set
To address RQ1’s concern with the shift in rule prevalence over time, we constructed a Longitudinal Subreddit Set, using a July 2023 data collection effort as a baseline, and identifying the subset of subreddits from the 2024 crawl that were seen in the earlier crawl. ChatGPT launched at the end of 2022, kicking off a period of increased public awareness and usage of AI. Given the dynamic nature of such a moment of flux, a longitudinal analysis could provide insight into the evolution of community norms and practices in a significant period soon after this launch.
@@ 3.3. Data Labeling
To answer our research questions about rules governing the use of AI, we created and applied labels to both the communities with rules and the AI rules contained in our sample. Through the process described below, we used a hybrid human-LLM coding technique to determine whether each of the subreddits in the Rules Subreddit Set and the Longitudinal Subreddit Set had an explicit rule about AI use, and to label them accordingly. We used these labels to identify the subset of subreddits with AI rules. We refer to the subset of the Rules Subreddit Set with AI rules as the AI Rules Subreddit Set ($N=4,251$). The specific AI-related rules from this set of subreddits became our AI Rules Set ($N=4,458$). We used an iterative inductive coding technique to identify common categories that emerged from the AI rules, and a hybrid human-LLM coding technique to label the rules according to these categories, which we describe in detail below.
Section 3.3.1 details our process for detecting AI rules. Next, Section 3.3.2 describes how we identified categories of AI rules. Section 3.3.3 then describes how we labeled our data with these categories, and then finally, Section 3.3.4 describes how we applied community type labels from several existing frameworks to all subreddits in the Rules Subreddit Set.
@@
3.3.1. Detecting Community Rules Governing AI Use
Community rules are often long and complex, which made our datasets too large to manually code for the presence of AI rules. To get a rough sense of the prevalence of these rules, we first used a simple regex to search the subreddits’ rules for AI-related keywords. This step found matches in the rules of less than $2\%$ of the subreddits. We then devised a protocol, inspired by recent HCI work using LLMs for qualitative labeling tasks (Choksi et al., 2024; Rao et al., 2024; Aubin Le Quéré et al., 2024), to detect AI-related rules across the larger datasets with more precision. The approach is based on comparing LLM-produced labels with those produced by human experts and iterating on the LLM prompt until desirable inter-rater reliability (IRR) is achieved between the human and LLM labels.
[…]
We used these manually labeled rules as ground truth against which to evaluate the performance of an LLM coder, which we planned to use to label the rest of the data. We crafted a prompt that combined the human coder instructions with Chain-of-Thought prompting techniques (Wei et al., 2023), including the recommended eight few-shot, in-context examples, which we selected from the manually labeled sample. Using OpenAI’s API, we produced LLM labels by prompting gpt-4o-mini-2024-07-18, OpenAI’s flagship small model at the time, with this prompt and the same input data that were given to the human coders (a subreddit’s rules). We compared the LLM-produced labels with the human-produced ones and calculated an IRR of .7.7.7.7, which we deemed insufficient. Two human coders examined the discrepancies, discussed what they thought the correct label should be, and updated the prompt with a clarifying example. For example, one rule labeled NO by humans but YES by the LLM was about bots: “No bots(No bots): No bots. All bots are banned unless the bot provides exceptional value to /r/ MealPrepSunday.” To better align the LLM label with our preferred definition of an AI rule, we added this rule as a negative example in the prompt, with the explanation:“This rule bans bots, not AI. Bots are not necessarily AI.”
To improve IRR we repeated this process of randomly sampling, manually coding, generating LLM labels, comparing the results, and updating the prompt with examples that the model got wrong. We performed three iterations of this step, at which point the LLM labels had an IRR of .95.95.95.95 with the human labels. We then generated labels for the entirety of both the Longitudinal Subreddit Set and the Rules Subreddit Set by prompting the gpt-4o-mini-2024-07-18 model via OpenAI’s API. The final prompt used can be found in Appendix A.1. This process identified the AI Rules Subreddit Set ($N=4,251$) of subreddits with rules governing AI use.
3.3.2. Identifying an AI Rule Taxonomy
In order to better understand the types of rules that subreddits use to govern AI, we identified these rules in our sample and used an iterative inductive coding process to develop a descriptive AI Rule Taxonomy. To identify these rules, we slightly modified the prompt developed and validated in Section 3.3.1 so that it would flag individual rules rather than a set of rules as being AI-related. Since the prompt required only a slight modification from the prior prompt, and since human coders would be manually reviewing the labels in the next step, we decided that we did not need to validate model performance before tagging the entire sample. We used this prompt to instruct the gpt-4o-mini-2024-07-18 model to produce labels for each of the $N=33,674$ individual rules in the AI Rules Subreddit Set, which yielded the AI Rules Dataset ($N=4,458$) of individual rules about AI.
Next, two of the authors collaborated to identify emergent labels for the rules. The authors randomly selected $N=25$ rules from this dataset and independently performed affinity mapping (Harboe and Huang, 2015) by grouping rules that appeared similar into emergent categories. The authors then compared results and consolidated their categories into labels that best captured the elements of interest. Each of the two authors then independently applied this set of labels to another random sample of $N=25$ rules, once again grouping similar rules, including into new emergent groups where the existing labels did not suffice. The two authors repeated this process of randomly sampling, independently labeling, collaboratively comparing their results, and consolidating their labels, until no new labels emerged and they were confident that their taxonomy covered the relevant dimensions of variation in the dataset. The final AI Rules Taxonomy contained $28$ labels grouped into four categories: Rationale, Stance, Medium and Framing, which are listed in Table 1 alongside example labels for each category (see Table 6 for the full taxonomy and examples rules for each label). We labeled the entire AI Rules Dataset according to this taxonomy, as we describe next.
@@ 3.3.3. Applying Rule Classification Labels
We used LLMs to label all rules in the AI Rules Dataset with labels from the AI Rules Taxonomy and labels from an additional Rule Type category provided by Fiesler et al. (Fiesler et al., 2018). The two first authors collaboratively developed a codebook with instructions for applying these labels. Fiesler et al. does not provide codebook instructions, so the authors interpreted the examples presented in the study and came up with instructions for how to apply each of the $24$ possible labels. Rule Type labels are not mutually exclusive, so we represented them as a series of binary labels, one for each Rule Type label.
@@ 3.3.4. Applying Subreddit Classification Labels
We also sought to label subreddits in the Rules Subreddit Set with details about their community type. To this end, we devised a process to apply labels from two label categories used in prior HCI literature: Topic labels from Fiesler et al. (Fiesler et al., 2018) and Community Archetype labels from Prinster et al. (Prinster et al., 2024). In applying labels, we considered a subreddit’s name, description, and rules metadata fields. The label categories for subreddits can be found in Table 2 alongside example labels for each category.
@@ 4. Findings
We group our results by research question and present in sequence our findings on: (1) the changes in the prevalence of community rules governing AI use, (2) the types of communities that have these rules, and (3) details about the rules themselves.
Table 3. Prevalence of rules and rules governing the use of AI in the Longitudinal Subreddit Set for (1) all subreddits and (2) the top 1% largest subreddits, ranked by subscriber count at the time of the 2024 crawl.
July 2023 November 2024
N % of sample N % of sample % Change
@@
4.1. RQ1: How Has the Prevalence of AI Rules Changed Over Time?
Figure 2. Number of subreddits in the Longitudinal Subreddit Set with AI rules. Subreddits are bucketed into deciles based on their subscriber count at the time of the second crawl. Labels above the bars indicate percent change between crawls.
@@ 4.1. RQ1: How Has the Prevalence of AI Rules Changed Over Time?
The image contains a bar chart showing the number of AI-related rules in subreddits during two different crawls, with subreddits bucketed into deciles by subscriber count. The Y-axis represents the number of subreddits and the X-axis is subscriber count decile at the time of the second crawl (measured by the lower edge of each decile). Each decile has two bars above it: the left bar represents the 2023 crawl and the right represents the 2024 crawl. The number of subreddits with AI rules steadily increases as subscriber count grows. The largest decile (7,241) has the highest number, above 1750 during the second crawl, while the smallest decile (10) has a number near 0.
Table 3 shows the results of the detection task described in Section 3.3.1 on our Longitudinal Subreddit Set. We find that during the 2023 crawl, $N=1,348$ subreddits had rules governing the use of AI (.$6\%.$). By the time of our 2024 crawl, this number increased by $108.3\%$ to $N=2,808$ ($1.2\%$). Rules in general, and specifically rules governing the use of AI, are more common in larger subreddits. For example, $8.4\%$ of the largest $1\%$ of subreddits by subscriber count had AI rules during our first crawl. This percentage increased to $17.1\%$ in our second crawl. Figure 2 provides a more detailed demonstration of the differences in the prevalence of AI rules and the rates of increase in such rules among subreddits with different numbers of subscribers. The figure shows the number of subreddits in the Longitudinal Subreddit Set with AI rules at the time of each crawl. In order to expose the variation across subreddits of different sizes, we bucket subreddits along the X-axis into deciles based on their number of subscribers at the time of the second crawl (the labels represent the lower edge of the decile). Each decile contains approximately 22,773 subreddits. For each decile we plot two bars representing the number of subreddits with AI rules during the first crawl (left bar, lighter blue) and the second crawl (right, darker blue). The number above the bars represents the percent change between these two crawls. For example, the rightmost pair of bars shows that in the largest decile of subreddits, about $750$ had AI rules during the first crawl, and that this number increased by $128\%$ to about $1,750$ at the second crawl. Remember that each decile contains the same total number of subreddits. The figure thus shows that subreddits with more subscribers are more likely to have AI rules, and that communities of all sizes saw increases in the prevalence of AI rules between our two crawls, though that increase was more prominent for larger communities.
For comparison, we additionally analyzed the prevalence of AI rules in subreddits that are part of the more recent Broad Subreddit Set and that were created after our first crawl and were thus excluded from our Longitudinal Subreddit Set. We identified $N=30,962$ English language subreddits in the Broad Subreddit Set that were created after the first crawl but were at least a month old during the second crawl. Of these newly created subreddits, $N=467$ had AI rules ($1.5\%$), which is higher than the percentage of subreddits with AI rules during the second crawl in the Longitudinal Subreddit Set ($1.2\%$). While this difference may be due to the fact that new subreddits have different characteristics, our separate analysis in the following section shows that new subreddits are heavily featured in the distribution of subreddits with AI rules in our data.
@@ 4.2. RQ2: What Types of Communities Have Rules Governing AI Use?
In this section, we examine the characteristics of communities that have rules governing the use of AI. To analyze these characteristics, we ran two sets of analysis, one on the Broad Subreddit Set and another on the Rules Subreddit Set. The results from our Broad Subreddit Set analysis show that larger subreddits are most likely to have explicit rules regulating AI, compared to subreddits with no rules, as well as those with rules but without AI rules. The results from our Rules Subreddit Set analysis show that among subreddits with any rules, subreddits with AI rules are more likely to be associated with the Art or Celebrity topics than subreddits without AI rules.
4.2.1. Broad Subreddit Set Analysis
[…]
To validate our findings about differences in the subscriber count and age distributions between subreddit groups, we next ran a multinomial logistic regression on this same set of subreddits. We used as input variables subscriber count [^6] We log transform subscriber count to account for the variable’s skewed distribution. and age (in years), in an attempt to control for any interaction between the two. The dependent variable took one of three mutually exclusive values based on the rules associated with the subreddit during our crawl: has no rules, has rules but no AI rule, and has AI rule. This model found significant correlations between both input variables and the type of rules present. Compared to subreddits with rules but no AI rule, an increase in subscriber count increased the odds of having AI rules ($\beta=0.32,p<0.01$) while an increase in age decreased the odds ($\beta=-0.035,p<0.01$). Still using as baseline the subreddits with rules but no AI rule, the model output shows that an increase in subscriber count decreased the odds of having no rules ($\beta=-0.41,p<0.01$) while an increase in age increased these odds ($\beta=0.11,p<0.01$). These results further support the finding from our distribution comparison that larger subreddits are more likely to have AI rules. These results also provide a more nuanced understanding of the relationship between subreddit age and the presence of AI rules. While the distribution comparison found that subreddits with AI rules have the largest median age, the regression results suggest that this is partially due to an interaction between subscriber count and age. When controlling for subscriber count, the regression results suggest that subreddits with AI rules are actually more likely to be newer than both those with no rules and those with rules but no AI rules.
4.2.2. Rules Subreddit Set Analysis
Table 4. Differences in the prevalence of labels for Rules Subreddit Set subreddits ($N=99,969$) with and without AI rules. Out of the 28 topic labels, we show the ten with the largest absolute differences.
Label Subreddits w/out AI Rule Subreddits w/ AI Rule Difference
Community Archetype Content generation 51.68% 73.75% 22.07***
Affiliation with an entity 57.47% 63.3% 5.83***
Topical question and answer 7.29% 6.52% -0.78
Learning and perspective broadening 20.52% 16.61% -3.91***
Social support 10.29% 4.38% -5.91***
Topic Celebrity 3.87% 18.33% 14.45***
Art 2.36% 12.99% 10.63***
Entertainment 19.01% 22.96% 3.95***
Politics 2.21% 0.54% -1.67***
Business and Organizations 2.71% 0.99% -1.73***
Games 6.97% 4.68% -2.29***
Local 4.22% 1.11% -3.11***
Hobby 8.07% 4.7% -3.36***
Humor 9.75% 3.32% -6.43***
Support 8.98% 2.52% -6.46***